OpenStudy (anonymous):

obtain dy/dx 4xy+cos(x/y)=y

OpenStudy (anonymous):

first for the term 4xy \[4(y+x \frac{ dy }{ dx })\]

OpenStudy (anonymous):

for the second term cos(x/y) \[-\sin \frac{ x }{ y }*\frac{ y-x \frac{ dy }{ dx } }{ y ^{2} }\]

OpenStudy (anonymous):

\[4y+4x \frac{ dy }{ dx }-\frac{ 1 }{ y }*\sin \frac{ x }{ y }+\frac{ x }{ y ^{2} }*\sin \frac{ x }{ y }*\frac{ dy }{ dx }=\frac{ dy }{ dx }\]

OpenStudy (anonymous):

collect the coefficients od dy/dx

OpenStudy (anonymous):

\[(4x+\frac{ x }{ y ^{2} }*\sin \frac{ x }{ y }-1)=-4y+\frac{ 1 }{ y }*\sin \frac{ x }{ y }\]

OpenStudy (anonymous):

sorry ,the left hand side is multiplied by dy/dx

OpenStudy (anonymous):

Oh okay lol